OpenStudy (anonymous):

Differentiate with respect to x. 2(2x+1)^3

mathslover (mathslover):

\[\large{\frac{d}{dx}(2(2x+1)^3)}\]

mathslover (mathslover):

\[\large{2\frac{d}{dx}(2x+1)^3}\]

mathslover (mathslover):

m i wrong @JayDS ?

mathslover (mathslover):

Now we have to use chain rule

mathslover (mathslover):

that states that: \[\large{\frac{d}{dx}((2x+1)^3) =\frac{du^3}{du}*\frac{du}{dx}}\] where u = 2x+1 and \(\frac{du^3}{du}=3u^2\)

OpenStudy (anonymous):

wait, I'm confused, I think I was taught another way to differentiate.

mathslover (mathslover):

so what we have now is : \[\large{2(3(2x+1)^2 * \frac{d}{dx}(2x+1)}\] \[\large{2(3(2x+1)^2*2\frac{d}{dx}x +0}\]

mathslover (mathslover):

\[\large{6*2(2x+1)^2=12(x+1)^2}\]

mathslover (mathslover):

where r u having problem in understanding this @JayDS ??

OpenStudy (anonymous):

all of it.

mathslover (mathslover):

hmn.. do you know chain rule?

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

2(2x+1)^3 => 2*3(2x+1)^2 * 2 => ...

mathslover (mathslover):

@shubham.bagrecha u took that problem as wrong.. it is 2 * (2x+1)^3 not {2(2x+1)}^3

mathslover (mathslover):

hmn jayDS you must know chain rule first

OpenStudy (anonymous):

=> 6*2(2x+1)^2

OpenStudy (anonymous):

@mathslover i have done that only

OpenStudy (anonymous):

ok, lets say one of the example they used of y=(3x + 1)^2 D(3x + 1)^2 = 2(3x+1)^2-1 D(3x + 1) = 2 (3x+1) (3) = 6 (3x+1) but why do we have to use it D(3x+1) and why is it not raised to the power of 2? I'm confused.

OpenStudy (anonymous):

@mathslover

mathslover (mathslover):

wait lemme explain u ..

mathslover (mathslover):

\[\large{\frac{d}{dx}(3x+1)^2 = \frac{d}{dx}(9x^2+6x+1) }\] \[\large{\frac{d}{dx}(9x^2)+\frac{d}{dx}(6x)+\frac{d}{dx}(1)}\] \[\large{9\frac{d}{dx}(x^2)+6\frac{d}{dx}(x) + 0}\] \[\large{9 * 2x + 6 (1) + 0 }\] \[\large{18x+6}\]

OpenStudy (anonymous):

wait, I found out it is the derivative of the inside so therefore is just 3x + 1 but why do we use it?

OpenStudy (anonymous):

and sorry for my stupid question but why is it just d/dx, is it the same as dy=dx?

mathslover (mathslover):

d/dx means that we are differentiating a term with restect to x

mathslover (mathslover):

*respect

OpenStudy (anonymous):

ok, got it, I believe I know how to do it now and I blame my teacher for providing some exercises from another book that contain this stuff that she hasn't even taught us.

OpenStudy (anonymous):

and thanks.

mathslover (mathslover):

hmn no need to blame the teacher dude we are here... that is because openstudy is made

OpenStudy (anonymous):

haha yeh, that's why I appreciate the admin, website and people who help me on it. thankyou.